# Emergent symmetries in prethermal phases of periodically driven quantum systems

### v1

@@ II.1 Dynamical Freezing
A precise understanding of this timescale, for Magnus expansion, is given in Ref. 45; it predicts that the prethermal timescale after which the Magnus expansion breaks down and thermalization sets in is ${\rm O}(\exp[c\omega_{D}])$, where $c$ is typically a number determined by system timescale (for the present system, $c\sim J^{-1}$). Thus at low or intermediate frequencies, where $\omega_{D}c\leq 1$, thermalization sets in within the first few drive cycle; however for $\omega_{D}c\gg 1$, the prethermal timescale is exponentially large. This is shown in the inset of Fig. 1(a) which plots the time-averaged value of $M_{z}$ at large $n$ as a function of $h_{1}$ at $\omega_{D}=\omega_{2}^{\ast}$; it can be clearly seen that the system crosses over to a freezing regime at large $\omega_{D}$.
@@
Numerical analysis for finite chains at very large drive amplitude with $\omega_{D}=\omega_{1}^{\ast}$ indicates that the freezing timescale is not reached within $10^{5}$ drive cycles. This makes the freezing almost exact from an experimental point of view since it may easily exceed system lifetime in an ultracold platform. Such an exponential dependence of the prethermal timescale on the drive frequency or amplitude in a similar setup has been numerically verified in Ref. 79; this shall be discussed in Sec. III in more details. We note here that analogous phenomenon has been theoretically studied in the context of driven chaotic quantum dots [54], driven tilted Bose-Hubbard model [55], non-equilibrium Ising phase transitions [56], and spin-half staggered Heisenberg model in a transverse field [57]. It has also been experimentally investigated using ultracold ${}^{87}{\rm Rb}$ atoms [59] and nuclear spins where measurements were carried out using nuclear magnetic resonance (NMR) techniques [52].
@@
In the non-Hermitian context ($\gamma\neq 0$), such an emergent symmetry influences the steady state behavior. To understand this we note that in these systems, the time taken to reach the steady state depends on $\gamma$; consequently, as seen in Refs. 60, 61, at large drive frequencies, the steady state may be reached within this prethermal timescale where the properties of $H_{F}^{(1)}$ controls the steady state behavior. A detailed study of correlation functions [60] and entanglement dynamics [61] clearly indicates that this is the case. Interestingly, the behavior of half-chain entanglement entropy, $S_{L/2}$, also changes due to such emergent conservation. As shown in Ref. 61, for small $\gamma$, $S_{L/2}\sim\alpha\ln L$, while for large $\gamma$, $S_{L/2}$ shows an area-law behavior and thus become independent of $L$. In between, there is an transition, termed as entanglement transition in Ref. 61, which occurs at critical $\gamma_{c}$ where $\alpha\to 0$. It has been shown in Ref. 61 that the entanglement for such a driven chain, exhibits a logarithmic dependence for $\gamma\leq\gamma_{c}$, where $\gamma_{c}\sim J_{0}(\mu)$; consequently, $\gamma_{c}$ vanishes at the special drive frequencies. Thus, at the special drive frequencies the entanglement always follow an area law; this can be seen from Fig. 1(b) where the coefficient $\alpha$ is plotted as a function of $\gamma$ and $\omega_{D}$. We note that freezing, with focus on information scrambling, has also been studied for periodically kicked non-Hermitian rotors [62, 89].
@@
Before ending this section, we note that for Hermitian systems in all the above-mentioned cases, the prethermal timescale over which freezing takes place is finite, although it is exponentially large in the large drive amplitude regime. In contrast, there been recent claims regarding exact freezing in Ref. 58; the exact nature of such freezing, if it exists, remains to be rigorously proven. Also, we note here that although a single parameter drive usually does not lead to exact freezing, carefully chosen protocol may lead to freezing in non-interacting free fermion systems [64]. For interacting systems, a two-tone protocol can lead to exact dynamical freezing in interacting many-body systems via realization of exact Floquet flat bands; such freezing necessarily requires two parameters driven with frequencies having a specific ratio and can not be understood from the perspective of an emergent symmetry [65, 66, 67].
@@ II.2 Dynamical localization
The phenomenon of dynamical localization constitutes freezing of transport in a periodically driven quantum system [68, 69, 70, 71, 72]. This phenomenon has been studied in the context of interacting models with on-site periodic potentials [71], driven hardcore bosons in their superfluid phases [72], periodically kicked three-band models [73], kicked Bose-Hubbard chain [74], interacting, driven fermions [75], periodically kicked rotors [76], and in driven two-level systems studied using a Green function approach [77]. Experimentally, there has been a recent report of observing many-body dynamical localization [78]. It has some similarity with dynamic freezing since it constitutes drive induced freezing that leads to localization. This phenomenon has been studied in many different settings; in what follows, we shall understand the effect of approximate emergent symmetry behind it by using a simple model, namely, interacting fermions on a 1D lattice in the presence of an time-dependent electric field [68]. The Hamiltonian of such a model is given by
@@
$+U_{0}\sum_{j}\hat{n}_{j\uparrow}\hat{n}_{j\downarrow}-eE(t)\sum_{j,\sigma=\uparrow,\downarrow}j\hat{n}_{j\sigma}$
@@
In the non-interacting limit where $U_{0}=0$, $H_{\rm ferm}$ can be most easily analyzed using the equation of motion method outlined in Ref. 100. The Heisenberg equation of motion for $c_{j}$ is given by
$i\partial_{t}c_{j,\sigma}$ $=$ $-J^{\prime}(c_{j-1,\sigma}+c_{j+1,\sigma})-eE(t)jc_{j,\sigma}$ (12)
To solve this one moves to a rotating frame, using a transformation $\tilde{c}_{j,\sigma}(t)=U(t,0)c_{j,\sigma}(t)$, where
@@
$=$ $\exp[-ie\sum_{j}j\hat{n}_{j\sigma}{\mathcal{A}}(t)/\hbar]$
where ${\mathcal{A}}(t)$ is the vector potential with $\dot{\mathcal{A}(t)}=-E(t)$. In the transformed frame, the equation of motion $\tilde{c}_{j\sigma}$ admits an easy solution
@@
where $\dot{f}=-2J^{\prime}\cos(k+{\mathcal{A}}(t))$, and $\tilde{c}^{k}_{\sigma}$ is the annihilation operator of a fermion in the rotated basis with spin $\sigma$ at momentum $k$. Using this solution, one can find that the fluctuation of fermion occupation $\langle n^{2}\rangle=\langle\sum_{j}\hat{n}_{j}^{2}\rangle$ at half-filling to be [68]
@@
where $x=eE_{0}/(\hbar\omega_{D})$ and we have omitted the spin index for simplicity since Eq. 15 holds for both spins. We note that at special frequencies given by $x=\beta_{m}$ where $J_{0}(\beta_{m})=0$, the motion of the fermions are bounded since $\langle n^{2}\rangle$ is a bounded function of $t$, as can be seen from expressions of $\mu(t)$ and $\nu(t)$. Moreover, at stroboscopic times, $t=2\pi p/\omega_{D}$, where $p$ is an integer, $\langle n^{2}\rangle=0$; at these times, the system, if driven at special frequencies, returns to its initial state. These features are shown in Fig. 2 and constitute dynamical localization of non-interacting fermions.
@@
The stability of such dynamical localization in the presence of interaction requires an analysis of the interacting Hamiltonian. Such an analysis, carried out in Ref. 68, is most easily done by introducing the concept of dressed fermions whose annihilation operator, given by $f_{j\sigma\ell}$, is related to those of the bare fermions by
$c_{j\sigma}(t)$ $=$ $e^{-i\epsilon t}\sum_{\ell}f_{j\sigma\ell}e^{-i\ell\omega_{D}t}$ (16)
Here $f_{j\sigma\ell}$ can be thought to annihilate a bound state of a fermion with $\ell$ photons. The density of such composite objects can be represented by density operators $\hat{n}_{j\sigma\ell}=\sum_{\ell^{\prime}}f_{j\sigma\ell^{\prime}}^{\dagger}f_{j\sigma\ell+\ell^{\prime}}$. In terms of these composite operators, the Floquet Hamiltonian of the system has been found to be [68]
@@
To understand the nature and role of the approximate emergent symmetry in this context, one notes that $H_{F}$ has two energy scales, $x_{1}=U_{0}/J^{\prime}$ and $x_{2}=\hbar\omega_{D}/J^{\prime}$. In the limit when $x_{2}$ is the largest energy scale, only terms with $\ell=\ell^{\prime}$ in $H_{F}$ contribute and setting the drive frequency to the $m^{\rm th}$ zero of $J_{0}$ (i.e, $eE_{0}/(\hbar\omega_{D})=\beta_{m}$), one finds that there is no hopping of undressed fermions. We note that here we are restricted to large drive amplitude limit since the ratio $x=eE_{0}/(\hbar\omega_{D})$ is fixed. The dressed fermions hopping is finite, but suppressed by a factor of $1/\omega_{D}$, similar to the suppression of $H_{F}^{(2)}$ in the context of dynamic freezing (Eq. 9). Ignoring the latter terms at high drive frequencies, one finds that up to a large prethermal timescale, the bare fermion density, $c_{j\sigma}^{\dagger}c_{j\sigma}$, at any site $j$ commutes with the first order Floquet Hamiltonian. This indicates localization of undressed fermions since all hopping processes involving undressed fermions are suppressed. This prethermal dynamical localization at stroboscopic times is a consequence of the approximate emergent symmetry which conserves local fermion density. However, it is to be noted that achieving full localization is impossible in this system at any finite drive frequencies due to interaction leading to the presence of mobile dressed fermions; this feature has been alluded to different context in the literature where effect of interaction on dynamical localization has been studied [69, 70, 71, 72].
III Prethermal HSF and Floquet scars
@@
In this section, we shall first, in Sec. III.1, discuss the realization of prethermal HSF in driven system and the role of an emergent approximate symmetry behind it. This will be followed by a discussion of prethermal realization of quantum many-body scars (QMBS) in Sec. III.2.
III.1 Prethermal HSF
@@
The phenomenon of strong HSF was first pointed out for fractonic models [101, 102] and interacting spin model on a chain [103] in the context of time-independent Hamiltonians. Since then, this phenomenon has been studied using several models in many different contexts [104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 123, 122, 124, 125, 126, 127]. In earlier studies [101, 102, 103, 104], such fragmentation was shown to be always accompanied by two global symmetries; in most cases these symmetries correspond to number and dipole conservation. The phenomenon of HSF constitutes fragmentation of the system Hilbert space into an exponentially large number of dynamically disconnected sectors in the computational (often taken to be Fock) basis. In other words, for these systems it is possible to construct dynamically disconnected distinct Krylov subspaces, ${\mathcal{K}}_{j}$, such that the total Hilbert space dimension ${\mathcal{D}}_{t}=\bigoplus{\mathcal{K}_{j}}$. Such spaces can be span by repeated application of $H$ of an initial state $|\psi_{j}\rangle$; $\mathcal{K}_{j}=\mathcal{\text{span}}\{\ket{\Psi_{j}},\text{H}\ket{\psi_{j}},\text{H}^{2}\ket{\psi_{j}},...\}$. For strong HSF, there are an exponentially large number of such Krylov subspaces. The Hilbert space dimension (HSD) of the largest fragment (or equivalently largest Krylov subspace), ${\mathcal{D}}_{L}$, turns out to be exponentially smaller compared to the total HSD ${\mathcal{D}}_{t}$: ${\mathcal{D}}_{L}/{\mathcal{D}}_{t}\sim\exp[-cL]$ where $L$ is the system size in units of lattice spacing and $c$ is a ${\rm O}(1)$ non-universal constant. This leads to an ${\rm O}(e^{L})$ fragments and distinguishes this phenomenon from separation of Hilbert space into dynamically disconnected symmetry sectors where the number of such sectors is at most ${\rm O}(L)$. Such a fragmentation naturally leads to loss of ergodicity and constitutes a strong violation of eigenstate thermalization hypothesis (ETH) [128, 129, 130]. This phenomenon is often termed as strong Hilbert space fragmentation; for the rest of this section, we shall omit the qualifier ”strong” for brevity.
@@
It turns out that there are models which exhibit HSF without the presence of two or more global conserved quantities; such conservations could be emergent within a fragmented sector [106, 109, 113, 114, 117, 121]. Moreover, fragmentation need not only occur in the classical Fock basis, but can also happen in an entangled basis. This phenomenon was first pointed out as secondary fragmentation in Ref. 121; it has been later dubbed as quantum fragmentation and discussed in several works [110, 111]. Most of the models discussed for studying HSF are one-dimensional [119]; however, more recently some higher-dimensional models exhibiting HSF has been analyzed in details [123, 122, 124]. The effect of the range of interaction on HSF has also been studied using Fredkin chains with longer range interactions [125]. The presence of fragmentation in non-hermitian and open quantum models has also been investigated [126, 127].
@@
The prethermal Floquet version of this phenomenon has been discussed in Refs. 79, 80, 81, 82. Here we shall closely follow the model used in Refs. 79, 80, namely, a driven spinless fermionic chain with nearest and next-neighbor density-density interaction and nearest-neighbor hopping. The driven Hamiltonian, introduced in Ref. 79, is given by $H(t)=H_{0}(t)+H_{1}$, where
@@
Here $c_{j}$ denotes the fermion annihilation operator for the site $j$ of the chain, $\hat{n}_{j}=c_{j}^{\dagger}c_{j}$ is the fermion density operator, $V_{0}+V(t)$ and $V_{2}$ are the strengths of nearest- and next-nearest neighbor interactions respectively. The drive is implemented using a square-pulse protocol: $V(t)=-(+)V_{1}$ for $t\leq(>)T/2$, where $T=2\pi/\omega_{D}$ is the time period of the drive and $\omega_{D}$ is the drive frequency. We note that $H(t)$ conserves fermion density: $[H(t),N]=0$ where $N=\sum_{j}\hat{n}_{j}/L$ and $L$ is chain length.
Figure 3: (a)Plot of half-chain entanglement entropy $S(nT)/S_{p}$ as a function of $n$ and $\hbar\omega_{D}=V_{1}/2$ and several $V_{1}$ starting from a Fock state belonging to a fragment with ${\rm HSD}=1008$ and $S_{p}^{f}=0.8S_{p}$ (red dotted line). $S$ saturates to $S_{p}(S_{p}^{f})$ at low(high) $V_{1}$. (b) Plot of $S(nT)/S_{p}$ for random initial Fock states saturating to different $\eta$. (c) Plot of $C_{L}(nT)$ for $n=5000$ as a function of $\hbar\omega_{D}/V_{1}$ showing non-zero long time value of $C_{L}$ at special frequencies where $\hbar\omega_{D}/V_{1}=1/2,1/4$. (d) Plot of $C_{L}(nT)$ for $n=5000$ as a function of $V_{1}$ at $\hbar\omega_{D}=V_{1}/2$ for several chain length $L$ showing a clear crossover to a prethermal HSF regime at large $V_{1}$. The inset shows the number of cycles, $n^{\rm th}$, required for $C_{L}(nT)$ to reach its Floquet ETH predicted infinite-temperature value for $L=16$ at $\hbar\omega_{D}=V_{1}/2$; $n^{\rm th}$ shows exponential scaling with $V_{1}$ at large $V_{1}$. All figures are obtained from Ref. 79.
In the large drive amplitude regime, the leading term of the Floquet Hamiltonian can be computed within FPT and yields [79, 80]
@@
$-J\sum_{j}[(1-\hat{A}_{j}^{2})+\alpha(\gamma_{1})\hat{A}_{j}^{2}]c_{j}^{\dagger}c_{j+1}+{\rm h.c.}$

### v2

@@ II.1 Dynamical Freezing
A precise understanding of this prethermal timescale, for Magnus expansion, is given in Ref. mori1, . It demonstrates that the prethermal timescale after which the Magnus expansion breaks down and thermalization sets in is ${\rm O}(\exp[c\omega_{D}])$, where $c$ is typically a number determined by system timescale (for the present system, $c\sim J^{-1}$). Thus at low or intermediate frequencies, where $\omega_{D}c\leq 1$, thermalization sets in within the first few drive cycle. In contrast, for $\omega_{D}c\gg 1$, the prethermal timescale is exponentially large. This is shown in the inset of the left panel of Fig. 1 which plots the time-averaged value of $M_{z}$ at large $n$ as a function of $h_{1}$ at $\omega_{D}=\omega_{2}^{\ast}$. The plot clearly shows that the system crosses over to a freezing regime at large $\omega_{D}$.
@@
Numerical analysis for finite chains at very large drive amplitude with $\omega_{D}=\omega_{1}^{\ast}$ indicates that the freezing timescale is not reached within $10^{5}$ drive cycles. This makes the freezing almost exact from an experimental point of view since it may easily exceed the system lifetime in an ultracold platform. Such an exponential dependence of the prethermal timescale on the drive frequency or amplitude in a similar setup has been numerically verified in Ref. sg1, ; this shall be discussed in Sec. III in more details. We note here that analogous phenomenon has been theoretically studied in the context of driven chaotic quantum dots deb1 , driven tilted Bose-Hubbard model uma1 , non-equilibrium Ising phase transitions camilo1 , and spin-half staggered Heisenberg model in a transverse field apal1 . It has also been experimentally investigated using ultracold ${}^{87}{\rm Rb}$ atoms koch1 and nuclear spins where measurements were carried out using nuclear magnetic resonance (NMR) techniques adas3 .
@@
In the non-Hermitian context ($\gamma\neq 0$), such an emergent symmetry influences the steady state behavior. To understand this we note that in these systems, the time taken to reach the steady state depends on $\gamma$; consequently, as seen in Refs. tb1, ; tb2, , at large drive frequencies, the steady state may be reached within this prethermal timescale where the properties of $H_{F}^{(1)}$ controls the steady state behavior. A detailed study of correlation functions tb1 and entanglement dynamics tb2 clearly indicates that this is the case. Interestingly, the behavior of half-chain entanglement entropy, $S_{L/2}$, also changes due to such emergent conservation. As shown in Ref. tb2, , for small $\gamma$, $S_{L/2}\sim\alpha\ln L$, while for large $\gamma$, $S_{L/2}$ shows an area-law behavior and thus become independent of $L$. In between, there is a transition, termed as entanglement transition in Ref. tb2, , which occurs at critical $\gamma_{c}$ where $\alpha\to 0$. We note that such a transition is characterized by the drive frequency in contrast to the more well-known measurement-rate induced entanglement transition studied, for example, in Refs. fisher1 ; nahum1 ; huse1 . It has been shown in Ref. tb2, that the entanglement for such a driven chain, exhibits a logarithmic dependence for $\gamma\leq\gamma_{c}$, where $\gamma_{c}\sim J_{0}(\mu)$; consequently, $\gamma_{c}$ vanishes at the special drive frequencies. Thus, at the special drive frequencies the entanglement always follow an area law; this can be seen from the right panel of Fig. 1 where the coefficient $\alpha$ is plotted as a function of $\gamma$ and $\omega_{D}$. We note that freezing, with focus on information scrambling, has also been studied for periodically kicked non-Hermitian rotors huo1 ; liu1 .
@@
Before ending this section, we note that for Hermitian systems in all the above-mentioned cases, the prethermal timescale over which freezing takes place is finite, although it is exponentially large in the large drive amplitude regime. In contrast, there have been recent claims regarding exact freezing in Ref. adasnew, ; the exact nature of such freezing, if it exists, remains to be rigorously proven. Also, we note here that although a single parameter drive usually does not lead to exact freezing, carefully chosen protocol may lead to freezing in non-interacting free fermion systems rudner1 . For interacting systems, a two-tone protocol can lead to exact dynamical freezing in interacting many-body systems via realization of exact Floquet flat bands; such freezing necessarily requires two parameters driven with frequencies having a specific ratio and can not be understood from the perspective of an emergent symmetry tb0 ; skar1 ; skar2 .
@@ II.2 Dynamical localization
The phenomenon of dynamical localization constitutes freezing of transport in a periodically driven quantum system dynloc1 ; dynloc2 ; dynloc3 ; dynloc4 ; dynloc5 . This phenomenon has been studied in the context of interacting models with on-site periodic potentials dynloc4 , driven hardcore bosons in their superfluid phases dynloc5 , periodically kicked three-band models tanay1 , driven Bose-Hubbard model fava1 , interacting, driven fermions rg1 , periodically kicked rotors galit1 , in driven two-level systems studied using a Green function approach martinez1 , and to implement coherent destruction of tunneling between particles on a lattice tun1 . Experimentally, there has been a recent report of observing many-body dynamical localization guoexp as well as several works on dynamical localization in ultracold atoms ari1 It has some similarity with dynamic freezing since it constitutes drive induced freezing that leads to localization.
As mentioned above, dynamical localization has been studied in many different contexts. In what follows, we shall understand the effect of approximate emergent symmetry behind it by using a simple model, namely, interacting fermions on a 1D lattice in the presence of a time-dependent electric field dynloc1 . The Hamiltonian of such a model is given by
@@
$+U_{0}\sum_{j}\hat{n}_{j\uparrow}\hat{n}_{j\downarrow}-eE(t)\sum_{j,\sigma=\uparrow,\downarrow}j\hat{n}_{j\sigma},$
@@
In the non-interacting limit where $U_{0}=0$, $H_{\rm ferm}$ can be most easily analyzed using the equation of motion method outlined in Ref. dynloc6, . In this limit, the Heisenberg equation of motion for $c_{j}$ is given by
$i\partial_{t}c_{j,\sigma}$ $=$ $-J^{\prime}(c_{j-1,\sigma}+c_{j+1,\sigma})-eE(t)jc_{j,\sigma}.$ (12)
To solve this equation, one moves to a rotating frame, using a transformation $\tilde{c}_{j,\sigma}(t)=U(t,0)c_{j,\sigma}(t)$, where
@@
$=$ $\exp[-ie\sum_{j}j\hat{n}_{j\sigma}{\mathcal{A}}(t)/\hbar],$
and ${\mathcal{A}}(t)$ is the vector potential with $\dot{\mathcal{A}}(t)=-E(t)$. In the transformed frame, the equation of motion $\tilde{c}_{j\sigma}$ admits an easy solution
@@
where $\dot{f}=-2J^{\prime}\cos(k+{\mathcal{A}}(t))$, and $\tilde{c}^{k}_{\sigma}$ is the annihilation operator of a fermion in the rotated basis with spin $\sigma$ at momentum $k$. Using this solution, one can find that the fluctuation of fermion occupation $\langle n^{2}\rangle=\langle\sum_{j}\hat{n}_{j}^{2}\rangle$ at half-filling to be dynloc1
@@
where $x=eE_{0}/(\hbar\omega_{D})$ and we have omitted the spin index for simplicity since Eq. 15 holds for both spins.
From Eq. 15, we note that at special frequencies given by $x=\beta_{m}$, where $J_{0}(\beta_{m})=0$, the motion of the fermions are bounded since $\langle n^{2}\rangle$ is a bounded function of $t$. This can be seen from expressions of $\mu(t)$ and $\nu(t)$ in Eq. 15. Moreover, at stroboscopic times $t=2\pi p/\omega_{D}$, where $p$ is an integer, $\langle n^{2}\rangle=0$. At these times, the system, if driven at special frequencies, returns to its initial state. These features are shown in Fig. 2 and constitute dynamical localization of non-interacting fermions.
@@
The stability of such dynamical localization in the presence of interaction requires an analysis of the interacting Hamiltonian. Such an analysis, carried out in Ref. dynloc1, , is most easily done by introducing the concept of dressed fermions. These dressed fermions are bound state of a bare fermion with $\ell$ photons. The annihilation operator for the dressed fermions is given by $f_{j\sigma\ell}$; it is related to $c_{j\sigma}$ by
$c_{j\sigma}(t)$ $=$ $e^{-i\epsilon t}\sum_{\ell}f_{j\sigma\ell}e^{-i\ell\omega_{D}t}.$ (16)
The density of such composite objects can be represented by density operators $\hat{n}_{j\sigma\ell}=\sum_{\ell^{\prime}}f_{j\sigma\ell^{\prime}}^{\dagger}f_{j\sigma\ell+\ell^{\prime}}$. In terms of these composite operators, the Floquet Hamiltonian of the system has been found to be dynloc1
@@
To understand the nature and role of the approximate emergent symmetry in this context, one notes that $H_{F}$ has two energy scales, $x_{1}=U_{0}/J^{\prime}$ and $x_{2}=\hbar\omega_{D}/J^{\prime}$. In the limit when $x_{2}$ is the largest energy scale, only terms with $\ell=\ell^{\prime}$ in $H_{F}$ contribute. In this regime, for a drive frequency which satisfies $eE_{0}/(\hbar\omega_{D})=\beta_{m}$, one finds that there is no hopping of undressed fermions. We note that here we are restricted to large drive amplitude limit since the ratio $eE_{0}/(\hbar\omega_{D})$ is fixed.
In contrast, the dressed fermions hopping is finite, but suppressed by a factor of $1/\omega_{D}$, similar to the suppression of $H_{F}^{(2)}$ in the context of dynamic freezing (Eq. 9). Ignoring the latter terms at high drive frequencies, one finds that up to a large prethermal timescale, the bare fermion density, $c_{j\sigma}^{\dagger}c_{j\sigma}$, at any site $j$ commutes with the first order Floquet Hamiltonian. This indicates localization of undressed fermions in the prethermal regime.
This prethermal dynamical localization at stroboscopic times is a consequence of the approximate emergent symmetry which conserves local fermion density. However, this localization is approximate at any finite drive frequency. This is due to the interaction between the fermions which makes it impossible to achieve full localizations and leads to the presence of mobile dressed fermions. This feature has been alluded to in different contexts in the literature where effect of interaction on dynamical localization have been studied dynloc2 ; dynloc3 ; dynloc4 ; dynloc5 .
III Prethermal Hilbert-space fragmentation and Floquet scars
@@
In this section, we shall first, in Sec. III.1, discuss the realization of prethermal Hilbert-space fragmentation in driven system and the role of an emergent approximate symmetry behind it. This will be followed by a discussion of prethermal realization of quantum many-body scars (QMBS) in Sec. III.2.
III.1 Prethermal Hilbert-space fragmentation
The phenomenon of strong Hilbert-space fragmentation was first pointed out for fractonic models hsf1 ; hsf2 and interacting spin model on a chain hsf3 in the context of time-independent Hamiltonians. Since then, this phenomenon has been studied using several models in many different contexts hsf4 ; hsf5 ; hsf6 ; hsf7 ; hsf8 ; hsf9 ; hsf10 ; hsf11 ; hsf12 ; hsf13 ; hsf14 ; hsf15 ; hsf16 ; hsf17 ; hsf18 ; hsfrev ; hsfexp ; as1 ; frank2 ; ac1 ; sid1 ; gia1 ; nh1 ; open1 . In earlier studies hsf1 ; hsf2 ; hsf3 ; hsf4 , such fragmentation was shown to be always accompanied by two global symmetries; in most cases these symmetries correspond to number and dipole conservation.
The phenomenon of Hilbert-space fragmentation constitutes fragmentation of the system Hilbert space into an exponentially large number of dynamically disconnected sectors in the computational (often taken to be Fock) basis. In other words, for these systems it is possible to construct dynamically disconnected distinct Krylov subspaces, ${\mathcal{K}}_{j}$, such that the total Hilbert space dimension ${\mathcal{D}}_{t}=\bigoplus{\mathcal{K}_{j}}$. Such spaces can be span by repeated application of $H$ of an initial state $|\psi_{j}\rangle$; $\mathcal{K}_{j}=\mathcal{\text{span}}\{\ket{\Psi_{j}},\text{H}\ket{\psi_{j}},\text{H}^{2}\ket{\psi_{j}},...\}$.
@@ III.1 Prethermal Hilbert-space fragmentation
For strong Hilbert-space fragmentation, there are an exponentially large number of such Krylov subspaces. The Hilbert space dimension (HSD) of the largest fragment (or equivalently largest Krylov subspace), ${\mathcal{D}}_{L}$, turns out to be exponentially smaller compared to the total HSD ${\mathcal{D}}_{t}$: ${\mathcal{D}}_{L}/{\mathcal{D}}_{t}\sim\exp[-cL]$, where $L$ is the system size in units of lattice spacing and $c$ is a ${\rm O}(1)$ non-universal constant. This leads to an ${\rm O}(e^{L})$ fragments and distinguishes this phenomenon from separation of the Hilbert space into dynamically disconnected symmetry sectors; the number of the latter sectors is at most ${\rm O}(L)$. Such a fragmentation naturally leads to loss of ergodicity and constitutes a strong violation of eigenstate thermalization hypothesis (ETH) ethref1 ; ethref2 ; ethref3 . This phenomenon is often termed as strong Hilbert space fragmentation; for the rest of this section, we shall omit the qualifier ”strong” for brevity.
@@
It turns out that there are models which exhibit Hilbert-space fragmentation without the presence of two or more global conserved quantities; such conservations could be emergent within a fragmented sector hsf6 ; hsf9 ; hsf13 ; hsf14 ; hsf17 ; as1 . Moreover, fragmentation need not only occur in the classical Fock basis, but can also happen in an entangled basis. This phenomenon was first pointed out as secondary fragmentation in Ref. as1, ; it has been later dubbed as quantum fragmentation and discussed in several works hsf10 ; hsf11 . Most of the models discussed for studying Hilbert-space fragmentation are one-dimensional hsfrev ; however, more recently some higher-dimensional models exhibiting Hilbert-space fragmentation has been analyzed in details frank2 ; ac1 ; sid1 . The effect of the range of interaction on Hilbert-space fragmentation has also been studied using Fredkin chains with longer range interactions gia1 . The presence of fragmentation in non-hermitian and open quantum models has also been investigated nh1 ; open1 .
@@
The prethermal Floquet version of this phenomenon has been discussed in Refs. sg1, ; sg2, ; xu1, ; zhang1, . Here we shall closely follow the model used in Refs. sg1, ; sg2, , namely, a driven spinless fermionic chain with nearest and next-neighbor density-density interaction and nearest-neighbor hopping. The driven Hamiltonian, introduced in Ref. sg1, , is given by $H(t)=H_{0}(t)+H_{1}$, where
@@
Here $c_{j}$ denotes the fermion annihilation operator for the site $j$ of the chain, $\hat{n}_{j}=c_{j}^{\dagger}c_{j}$ is the fermion density operator, $V_{0}+V(t)$ and $V_{2}$ are the strengths of nearest- and next-nearest neighbor interactions respectively.
The drive is implemented using a square-pulse protocol: $V(t)=-(+)V_{1}$ for $t\leq(>)T/2$, where $T=2\pi/\omega_{D}$ is the time period of the drive and $\omega_{D}$ is the drive frequency. We note that $H(t)$ conserves fermion density: $[H(t),N]=0$ where $N=\sum_{j}\hat{n}_{j}/L$ and $L$ is chain length.
Figure 3: Top left: Plot of half-chain entanglement entropy $S(nT)/S_{p}$ as a function of $n$ and $\hbar\omega_{D}=V_{1}/2$ and several $V_{1}$ starting from a Fock state belonging to a fragment with ${\rm HSD}=1008$ and $S_{p}^{f}=0.8S_{p}$ (red dotted line). $S$ saturates to $S_{p}(S_{p}^{f})$ at low(high) $V_{1}$. Top right: Plot of $S(nT)/S_{p}$ for random initial Fock states saturating to different $\eta$. Bottom left: Plot of $C_{L}(nT)$ for $n=5000$ as a function of $\hbar\omega_{D}/V_{1}$ showing non-zero long time value of $C_{L}$ at special frequencies where $\hbar\omega_{D}/V_{1}=1/2,1/4$. Bottom Right: Plot of $C_{L}(nT)$ for $n=5000$ as a function of $V_{1}$ at $\hbar\omega_{D}=V_{1}/2$ for several chain length $L$ showing a clear crossover to a prethermal Hilbert-space fragmentation regime at large $V_{1}$. The inset shows the number of cycles, $n^{\rm th}$, required for $C_{L}(nT)$ to reach its Floquet ETH predicted infinite-temperature value for $L=16$ at $\hbar\omega_{D}=V_{1}/2$; $n^{\rm th}$ shows exponential scaling with $V_{1}$ at large $V_{1}$. All figures are obtained from Ref. sg1, .
In the large drive amplitude regime, the leading term of the Floquet Hamiltonian can be computed within FPT and yields sg1 ; sg2
@@
$-J\sum_{j}[(1-\hat{A}_{j}^{2})+\alpha(\gamma_{1})\hat{A}_{j}^{2}]c_{j}^{\dagger}c_{j+1}+{\rm h.c.},$
where $\hat{A}_{j}=(\hat{n}_{j+2}-\hat{n}_{j-1})$, $\gamma_{1}=V_{1}T/(4\hbar)$ and $\alpha(\gamma_{1})=\gamma_{1}^{-1}\sin\gamma_{1}\exp[i\gamma_{1}\hat{A}_{j}]$.